OpenStudy (anonymous):

show that, y^2 + 2e^(-ky) +6 given y" at point (0,2) is -3/2

OpenStudy (dan815):

i dont understand question can u clarify

hartnn (hartnn):

show what ?

hartnn (hartnn):

or u need to find k, i think..

OpenStudy (anonymous):

opps, sorry, the equation shuold be y^2 + 2e^(-ky) = 6

hartnn (hartnn):

still, it isn't clear.... you have an equation, y^2 + 2e^(-ky) = 6 and given y" at point (0,2) is -3/2 what you need to show or find ?

OpenStudy (anonymous):

sorry for the trouble. the qn i got is blur... The qn is show that, for y^2 + 2e^(-xy) = 6, the value of y" at the point (0,2) is - 3/2

OpenStudy (anonymous):

$$y^2+2e^{-xy}=6\\\frac{d}{dx}\left[y^2+2e^{-xy}\right]=\frac{d}{dx}[6]\\2y\frac{dy}{dx}-2\left(y+x\frac{dy}{dx}\right)e^{-xy}=0\\2\frac{dy}{dx}\left(y-xe^{-xy}\right)=2ye^{-xy}\\\frac{dy}{dx}=\frac{ye^{-xy}}{y-xe^{-xy}}=\frac{y}{ye^{xy}-x}\\\quad\implies\frac{dy}{dx}=\frac2{2e^0-0}=1\\\frac{d^2y}{dx^2}=\frac{ye^{xy}-x-y(ye^{xy}(y+x\frac{dy}{dx})+\frac{dy}{dx}e^{xy}-1)}{(ye^{xy}-x)^2}\\\quad\implies\frac{d^2y}{dx^2}=\frac{2e^0-0-2(2e^0(2+0)+e^0-1)}{(2e^0-0)^2}=\frac{2-2(4)}{4}=\frac{2(-3)}{4}=-\frac32$$

OpenStudy (dan815):

wow thats some crazy latex work

OpenStudy (anonymous):

Thanks a lot! =D